# Babel-ImageNet: Massively Multilingual Evaluation of Vision-and-Language Representations

### v1

@@ Appendix E Full Results
lang OAI B32 ST B32 MC mB B32 OC B32 MC B32 MC B16+ MC L14 AC L14 OC H14
en 61.3 38.2 29.2 62.8 42.6 46.4 51.6 69.9 77.1
af 10.7 14.8 24.9 36.7 46.9 49.9 55.2 22.2 47.4
am 1.5 0.9 1.4 1.0 28.8 27.1 29.0 7.6 2.7
ar 0.5 10.6 13.8 24.6 30.1 31.5 35.9 41.7 31.9
as 1.0 2.9 17.5 5.4 22.6 21.4 24.4 6.5 9.8
az 7.0 12.6 23.2 26.0 25.2 26.1 28.6 16.3 37.1
be 0.7 9.9 14.1 27.1 32.1 34.6 36.4 27.1 40.2
bg 0.8 18.9 21.1 41.5 39.5 42.3 46.8 34.0 54.1
bn 0.2 3.0 25.7 8.7 29.1 31.1 35.1 7.1 19.6
br 7.8 7.5 7.8 12.8 12.0 11.8 13.2 12.3 16.0
bs 12.9 35.3 38.9 56.6 58.9 59.5 65.2 29.4 70.1
ca 11.2 17.2 19.5 33.3 33.9 36.9 39.6 31.4 46.8
cs 6.3 19.2 21.6 42.1 35.7 38.0 41.5 19.6 54.1
cy 4.9 5.5 19.4 12.0 10.9 11.1 12.6 12.6 16.4
da 12.1 22.4 22.3 43.4 43.3 46.6 51.0 24.7 55.5
de 15.3 18.3 19.4 47.8 38.7 41.7 45.6 27.6 61.2
el 0.9 15.1 16.3 37.2 38.5 41.2 45.2 4.9 51.0
eo 5.8 10.7 12.9 21.9 24.0 24.7 28.0 24.7 28.7
es 16.6 20.4 19.7 45.2 35.4 38.2 41.4 51.2 57.6
et 4.5 13.3 19.6 29.8 41.4 44.7 48.1 15.5 37.8
eu 7.5 8.7 9.5 17.1 18.3 18.7 21.0 17.6 21.4
fa 0.4 13.6 20.3 32.7 25.0 27.8 28.7 20.1 42.6
fi 3.8 11.1 14.1 25.2 25.1 27.1 29.7 10.4 35.7
fr 20.7 20.6 19.8 46.6 35.0 38.4 41.7 53.6 59.9
fy 12.8 17.8 17.5 35.9 35.8 35.6 37.0 25.8 42.1
ga 2.6 2.1 3.0 6.6 8.2 9.1 9.4 8.3 9.5
gd 3.0 2.7 3.4 6.0 7.3 8.4 10.4 6.1 7.4
gl 15.3 23.6 22.5 44.3 39.6 40.9 44.5 43.4 54.6
gu 0.4 7.4 28.5 6.8 26.7 27.8 32.4 11.2 12.0
ha 1.5 1.8 19.8 2.2 4.5 4.6 5.3 3.3 3.6
he 0.3 12.7 16.5 26.7 20.7 21.9 23.5 7.0 35.9
hi 0.1 17.3 20.9 25.8 36.4 37.0 41.1 12.8 38.9
hr 8.6 25.2 29.7 45.1 50.0 50.6 56.7 21.4 58.3
hu 5.4 16.8 19.8 41.1 41.8 44.9 49.8 16.4 53.7
hy 0.2 7.3 16.5 9.9 17.6 18.2 18.2 5.3 17.5
id 15.7 23.1 26.2 44.2 46.8 50.0 54.6 26.2 57.6
is 2.9 6.1 18.7 13.6 39.3 41.4 45.3 10.6 19.3
it 15.1 19.3 19.6 42.1 34.7 37.6 40.6 48.6 55.1
ja 4.1 15.5 18.1 42.7 24.6 25.0 28.7 55.6 57.3
jv 15.0 15.0 19.2 31.8 32.6 33.4 37.1 21.4 40.6
ka 0.2 8.5 15.1 11.0 21.6 22.5 23.8 8.9 20.4
kk 0.6 10.7 19.7 28.1 25.5 26.2 27.9 31.5 34.8
km 0.9 0.7 0.6 3.3 17.5 16.5 19.3 6.7 3.3
kn 1.0 2.1 21.1 4.0 27.2 27.2 29.2 10.8 5.7
ko 0.4 12.2 15.9 33.8 21.4 23.7 24.7 53.2 43.7
ku 5.1 8.9 9.4 14.2 14.1 12.5 14.6 11.2 16.0
ky 0.6 13.9 18.3 32.3 28.3 28.3 31.4 32.1 38.6
la 12.3 12.5 10.5 22.2 20.2 19.8 22.2 23.9 28.0
lo 0.7 0.7 0.7 0.7 12.5 12.4 13.0 9.9 2.1
lt 5.9 18.2 21.4 35.2 27.5 29.4 31.4 18.5 45.7
lv 7.8 23.4 25.7 38.8 33.5 33.4 36.7 21.3 47.2
mg 4.1 9.0 8.6 14.0 14.2 13.0 14.0 13.5 14.8
mk 1.3 20.0 23.3 36.0 45.6 47.2 50.9 31.1 49.6
ml 0.2 1.3 19.9 1.7 38.3 38.3 43.5 7.6 4.5
mn 1.0 13.9 28.0 18.7 19.4 20.9 21.5 20.2 26.9
mr 1.2 29.2 25.7 32.4 45.1 45.2 47.5 17.4 41.9
ms 14.4 18.8 21.6 37.4 36.6 38.6 43.1 20.9 48.5
my 0.5 10.2 2.4 5.8 14.6 14.5 16.0 9.4 10.4
ne 0.6 16.6 17.2 25.3 39.5 37.5 39.2 16.3 36.5
nl 12.7 18.5 18.2 42.0 36.5 38.6 43.9 24.7 55.1
no 9.2 19.2 20.7 42.1 39.7 43.0 46.9 23.1 53.7
om 5.0 8.9 3.8 5.7 8.8 11.9 10.0 8.2 13.3
or 1.9 1.7 1.7 3.2 31.0 30.8 35.9 12.4 1.7
pa 1.5 4.0 8.0 1.8 32.2 33.7 31.5 11.5 2.8
pl 7.4 15.9 17.9 39.3 32.6 35.4 38.4 19.6 51.4
ps 1.3 11.6 11.1 21.7 18.8 19.4 23.0 17.9 24.9
pt 14.1 22.6 21.1 47.6 39.5 42.8 46.8 37.1 59.7
ro 9.9 15.9 18.5 35.2 36.8 39.9 43.7 23.9 47.8
ru 0.6 15.8 18.0 43.2 34.5 36.3 41.3 48.3 58.4
sa 0.6 9.5 9.4 10.9 17.0 17.3 18.8 11.0 11.5
sd 1.7 8.6 20.8 15.0 27.9 27.6 29.6 16.0 17.0
si 2.7 2.1 2.2 3.1 33.6 33.4 36.2 18.7 5.6
sk 8.5 22.4 24.9 45.1 39.7 42.5 45.5 22.7 58.8
sl 5.9 22.0 24.2 36.9 42.3 44.4 48.4 17.2 48.8
so 2.4 5.0 17.6 6.3 12.7 10.1 12.4 3.9 7.4
sq 8.1 24.3 25.5 33.9 47.1 49.3 53.6 21.6 43.6
sr 1.2 17.1 18.9 33.8 43.7 45.5 49.5 30.4 47.8
su 12.1 11.6 15.9 27.9 29.2 27.8 29.2 18.6 30.6
sv 8.9 19.2 20.6 43.1 41.2 45.5 48.9 20.9 55.3
sw 4.7 6.4 17.0 10.1 36.2 36.5 38.7 10.3 13.0
ta 0.3 2.2 18.4 4.5 18.0 18.7 20.6 5.5 6.7
te 0.5 3.1 24.7 2.7 31.0 32.4 34.1 10.8 3.5
th 1.3 10.5 15.3 28.7 27.2 29.3 32.4 11.7 40.2
tl 7.8 8.2 15.8 17.6 31.9 33.0 37.6 18.0 21.8
tr 7.5 19.2 22.6 41.5 41.7 44.5 47.9 17.4 53.0
ug 1.3 2.3 2.1 3.2 11.5 12.7 12.5 9.0 4.8
uk 0.6 14.9 17.3 39.2 35.1 36.1 40.5 34.4 53.4
ur 0.6 17.5 26.8 25.8 29.4 26.9 30.8 18.6 37.2
uz 6.2 9.7 21.5 21.4 22.0 21.7 23.7 17.1 28.2
vi 6.8 17.5 18.7 41.0 37.0 39.2 43.7 11.5 53.4
xh 18.7 14.2 15.8 24.4 17.7 19.1 20.2 22.2 27.7
yi 0.6 1.3 1.0 2.5 18.7 18.1 19.3 5.3 4.9
zh 1.8 19.3 21.7 40.9 32.7 36.0 40.4 52.7 53.5
model en de es id ja ru tr zh average
OpenAI B32 51.8 10.7 20.9 4.1 1.7 0.5 2.1 0.5 5.8
OpenCLIP XLMR B32 62.9 53.6 61.6 48.0 48.6 63.3 52.2 53.2 54.3
OpenCLIP XLMR H14 73.9 67.6 77.0 65.6 65.2 77.6 68.7 70.0 70.2
M-CLIP XLMR B32 53.3 47.1 52.8 48.1 39.1 55.2 51.0 49.2 48.9
M-CLIP XLMR B16+ 65.3 60.4 66.7 61.0 49.7 70.6 65.0 61.2 62.1
M-CLIP XLMR L14 60.6 52.5 60.3 55.3 44.6 61.3 58.1 53.5 55.1
AltCLIP XLMR L14 64.6 32.5 64.8 19.8 55.1 65.5 9.7 60.3 43.9
M-CLIP mBERT B32 47.0 38.1 45.5 38.7 36.4 44.6 40.8 41.9 40.8
ST mBERT B32 46.3 31.7 39.6 31.1 28.1 35.2 29.4 34.7 32.8
@@
lang OAI B32 OC B32 OC H14 MC B32 MC B16+ MC L14 AC L14 MC mB B32 ST B32
en 40.4 49.9 55.5 39.6 49.5 40.9 43.9 32.4 32.7
average 4.2 43.2 51.4 41.5 52.1 44.5 21.6 29.1 20.6
ar 0.2 38.9 48.3 41.6 51.9 43.9 43.8 24.7 17.0
bn 0.0 2.1 5.5 25.1 36.5 26.3 1.3 15.1 0.2
cs 1.8 47.3 56.0 42.0 53.4 44.2 10.0 29.9 23.1
da 3.5 53.1 63.6 50.6 63.2 55.1 12.1 34.3 27.0
de 9.8 64.5 73.2 53.8 67.8 57.6 30.0 37.6 30.1
el 0.1 45.9 55.5 43.3 52.9 46.7 3.9 26.9 18.8
es 15.4 54.6 62.8 46.4 56.3 49.1 49.6 33.3 27.3
fa 0.1 49.4 58.2 41.4 52.5 44.0 13.1 31.9 21.7
fi 1.6 44.6 58.1 47.1 58.8 49.8 5.7 32.6 22.3
fil 3.7 7.2 9.8 36.6 47.3 40.0 7.3 22.9 3.7
fr 19.5 61.2 70.5 51.2 63.9 55.5 55.8 37.9 31.9
he 0.2 48.9 58.5 38.3 50.2 42.8 6.3 32.2 19.1
hi 0.1 17.5 22.3 20.6 30.7 23.3 3.1 11.7 8.2
hr 1.7 49.8 61.1 49.9 62.0 53.1 8.5 36.8 27.3
hu 1.8 50.4 63.8 51.8 63.6 55.9 9.4 30.6 23.1
id 5.7 56.7 67.1 53.2 65.9 57.4 18.8 38.3 30.8
it 8.0 59.3 68.6 50.5 63.0 53.4 51.9 35.7 26.6
ja 1.7 59.7 69.4 42.8 54.1 46.6 58.3 35.3 27.0
ko 0.1 44.5 53.7 36.6 46.7 41.1 51.7 27.5 17.6
mi 0.2 0.5 0.4 0.2 0.3 0.3 0.3 0.1 0.2
nl 8.0 49.3 57.3 44.8 54.8 47.6 21.2 31.2 25.1
no 3.5 50.1 61.0 48.8 60.4 52.1 12.7 34.9 24.0
pl 1.7 55.8 65.5 49.0 59.9 53.3 13.5 33.3 26.8
pt 10.1 55.5 65.4 46.5 57.8 51.0 37.7 32.4 28.3
quz 1.2 3.0 3.5 1.8 2.0 1.9 3.0 1.3 1.2
ro 3.4 57.4 68.2 51.9 65.9 55.9 16.5 33.6 25.2
ru 0.5 64.6 72.7 53.0 65.6 56.6 56.1 38.1 30.1
sv 3.0 52.9 61.5 49.4 60.4 53.1 12.7 34.0 25.3
sw 1.1 2.3 3.2 30.0 39.3 32.6 2.5 17.5 0.9
te 0.0 0.3 0.6 20.1 27.6 21.2 4.1 18.3 0.0
th 1.3 45.3 55.4 42.2 54.0 44.7 13.3 28.1 15.8
tr 1.4 47.2 58.0 45.0 56.1 48.9 7.2 29.9 21.0
uk 0.2 56.2 65.9 51.8 62.8 54.9 33.3 35.6 26.8
vi 0.8 54.9 65.4 48.8 60.9 51.3 5.6 33.9 26.6
zh 0.3 56.4 62.9 46.4 58.3 50.1 54.8 36.5 28.3
@@
OpenAI B32 54.6 12.9 18.8 22.1 11.5 2.1 0.2 2.7 0.7 1.6 1.2 7.4
OpenCLIP XLMR B32 63.0 54.4 56.7 54.9 55.1 45.8 44.3 55.0 50.5 50.0 51.0 51.8
OpenCLIP XLMR H14 72.4 66.1 66.8 66.4 64.1 62.1 56.1 69.2 61.8 63.3 63.8 64.0
M-CLIP XLMR B32 54.7 50.2 49.0 50.9 49.0 38.4 41.5 52.7 47.3 50.3 49.1 47.8
M-CLIP XLMR B16+ 66.6 63.9 62.8 63.5 62.9 49.6 53.5 63.9 57.1 62.7 60.6 60.1
M-CLIP XLMR L14 59.1 54.6 56.8 56.5 56.4 43.2 45.1 57.3 51.7 55.6 55.6 53.3
AltCLIP XLMR L14 62.9 35.6 58.8 60.5 56.8 52.2 55.6 18.0 53.6 11.0 59.7 46.2
M-CLIP mBERT B32 50.4 45.0 45.3 45.8 45.0 38.2 41.7 47.1 40.1 44.6 45.2 43.8
ST mBERT B32 49.2 36.8 38.2 41.0 38.5 28.7 25.9 36.1 32.4 33.2 37.6 34.8

### v2

@@ Appendix D Full Results
model en de es id jp ru tr zh average
AltCLIP XLMR-L L-14 64.50 33.05 66.65 20.75 57.05 66.00 10.65 62.50 45.24
M-CLIP mBERT B-32 42.45 32.20 39.80 33.15 31.10 39.10 36.05 36.60 35.43
M-CLIP XLMR-L B-16+ 63.80 59.10 67.35 62.75 50.45 72.35 66.25 63.20 63.06
M-CLIP XLMR-L B-32 49.15 42.70 48.55 43.85 33.50 51.60 47.00 44.65 44.55
M-CLIP XLMR-L L-14 58.00 50.85 58.45 52.95 42.40 59.00 55.85 52.10 53.09
NLLB-SigLIP-base 66.80 58.30 67.45 60.65 53.10 70.65 64.90 59.25 62.04
NLLB-SigLIP-large 70.55 65.00 72.30 65.10 60.20 74.55 68.90 63.20 67.04
OpenCLIP XLMR B-32 61.80 53.15 61.45 48.90 47.90 65.50 53.00 54.95 54.98
OpenCLIP XLMR-L H-14 73.85 66.85 77.65 64.80 63.70 78.60 68.85 69.90 70.05
mSigLIP 68.45 59.15 70.45 57.50 29.45 71.75 55.35 47.80 55.92
ST mBERT B-32 39.85 24.35 29.95 26.65 20.40 28.15 22.50 26.20 25.46
@@
lang AC L-14 MC mB B-32 MC B-16+ MC B-32 MC L-14 NS-base NS-large OC B-32 OC H-14 mSigLIP ST B-32
en 43.49 24.63 47.47 31.69 36.56 48.25 50.14 48.38 54.43 52.17 25.58
average 21.65 21.76 50.98 33.79 39.69 52.07 56.03 42.6 50.52 44.55 13.31
ar 43.56 18.03 51.12 34.34 38.93 53.02 55.0 39.26 47.73 45.13 9.75
bn 1.25 10.53 34.53 20.92 22.61 49.11 50.58 2.22 5.22 20.75 0.19
cs 9.8 21.76 50.23 33.63 38.13 49.81 53.78 45.08 53.41 46.95 14.57
da 12.0 27.11 61.85 42.86 50.56 58.41 63.35 52.29 62.73 54.32 18.57
de 30.17 28.69 65.79 45.13 53.2 63.79 68.32 63.68 72.44 67.06 19.62
el 4.08 19.64 51.76 35.51 41.45 49.07 55.02 45.0 54.08 40.44 12.49
es 48.75 24.24 55.36 37.6 43.55 54.41 58.22 54.02 61.52 59.53 17.49
fa 13.04 23.3 50.89 32.71 38.0 53.04 56.65 48.31 56.69 51.36 13.02
fi 5.79 24.16 58.1 38.04 44.02 57.68 63.25 42.95 57.15 45.03 13.43
fil 7.29 16.92 45.98 29.26 34.87 43.71 47.9 7.27 9.54 20.95 2.24
fr 55.4 28.4 62.51 42.2 50.36 62.24 66.63 60.79 69.72 64.53 21.3
he 6.57 23.9 48.75 30.39 37.82 56.97 61.44 48.17 57.93 51.01 10.99
hi 2.76 8.39 29.9 15.01 18.52 34.8 36.21 17.12 21.8 17.85 4.16
hr 8.53 29.31 61.41 41.96 48.72 57.51 62.01 48.82 59.68 48.63 18.97
hu 9.95 23.31 62.83 43.35 50.98 56.93 63.29 49.74 62.71 54.88 14.3
id 18.52 28.91 65.0 44.88 52.13 64.12 67.3 56.2 66.29 61.03 21.5
it 51.82 26.33 61.35 41.25 48.28 59.26 64.18 58.92 67.19 63.49 16.49
ja 58.65 27.1 54.15 33.83 40.38 58.29 61.64 59.14 69.55 42.66 16.19
ko 51.32 20.17 46.07 29.28 35.83 52.08 56.59 44.51 53.44 47.58 9.74
mi 0.23 0.02 0.34 0.13 0.27 23.99 27.68 0.51 0.49 0.32 0.11
nl 21.04 23.59 53.11 36.25 42.42 51.74 56.6 48.36 56.37 52.8 16.93
no 12.91 26.36 58.1 39.46 46.32 57.35 61.93 49.98 59.68 50.95 15.53
pl 13.68 26.26 59.03 41.27 48.45 56.62 62.74 55.31 65.2 58.03 17.98
pt 38.08 25.18 57.08 38.98 46.62 55.65 60.97 55.2 64.46 60.15 20.02
quz 2.69 0.81 2.14 1.19 1.61 16.63 18.6 2.93 3.5 3.03 0.75
ro 17.35 25.19 64.59 42.82 51.38 59.85 65.96 56.32 66.59 53.74 16.64
ru 56.54 28.83 65.17 44.24 51.92 61.83 65.75 63.61 72.36 68.18 20.17
sv 12.83 25.38 59.29 40.71 47.63 55.29 60.36 51.57 59.66 53.13 16.44
sw 2.5 12.52 39.02 22.75 27.59 41.98 44.78 2.5 3.18 11.1 0.41
te 3.86 13.18 26.83 14.54 17.24 37.71 39.99 0.25 0.58 4.35 0.06
th 13.38 20.43 52.29 33.21 38.72 53.87 56.49 44.44 54.29 25.01 9.38
tr 7.51 22.98 55.25 37.65 44.37 54.27 57.83 46.41 57.68 49.79 13.11
uk 33.67 26.32 61.75 42.83 49.6 57.44 61.98 55.37 65.18 54.87 17.24
vi 5.52 24.9 59.41 38.84 45.59 55.99 59.09 53.56 64.2 52.67 17.25
zh 54.74 26.44 56.97 37.59 44.05 51.98 54.96 55.34 61.92 50.32 16.67
@@
AltCLIP XLMR-L L-14 64.4 36.0 58.6 60.1 57.8 53.3 56.7 17.7 53.9 10.7 58.6 46.34
M-CLIP mBERT B-32 44.5 40.9 41.4 42.0 41.5 33.4 35.5 41.4 35.4 39.0 40.7 39.12
M-CLIP XLMR-L B-16+ 63.2 61.4 59.8 59.3 61.0 48.3 49.8 64.0 54.8 59.6 58.8 57.68
M-CLIP XLMR-L B-32 48.5 46.9 46.4 46.1 45.8 35.0 36.9 48.0 43.2 45.7 45.4 43.94
M-CLIP XLMR-L L-14 56.3 52.2 52.7 51.8 53.6 41.5 42.5 54.1 48.4 52.7 53.5 50.30
NLLB-SigLIP-base 70.8 64.2 66.3 66.0 66.2 55.3 61.2 68.0 61.6 66.0 60.2 63.50
NLLB-SigLIP-large 71.9 67.0 68.9 68.0 67.8 58.1 63.4 68.3 62.0 68.8 63.7 65.60
OpenCLIP XLMR B-32 63.2 54.5 54.6 55.7 55.7 47.1 43.8 55.5 50.3 48.2 50.8 51.62
OpenCLIP XLMR-L H-14 73.5 64.8 65.9 64.7 64.9 64.3 56.4 68.7 62.4 62.7 61.9 63.67
mSigLIP 68.0 60.8 62.7 59.9 58.1 32.9 49.6 59.6 56.9 55.9 50.4 54.68
ST mBERT B-32 43.3 32.3 32.9 32.0 29.8 21.4 18.5 28.8 25.5 25.3 31.0 27.75